Operations Report – November 20th

Crew 286 Operations Report 20-11-2023

SOL: 8

Name of person filing report: Donald "Duckie" Jacques

Non-nominal systems: Toilet Room door knob

Notes on non-nominal systems: It is becoming increasingly difficult to open the door. Upon inspection, I noted deep grooves in the actuator shaft for the latch, indicating the assembly is approaching end-of-life. The handle is still barely functional, however, sometimes it takes multiple attempts to open the door for exit.


Spirit rover used: No

Hours: (before EVA)

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging: Yes

Opportunity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Perseverance rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

General notes on rovers:

Summary of Hab operations: put notes here

WATER USE: 13.07 gallons

Water (static tank): 223.250 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes, in early morning

Summary of internet: Nominal

Summary of suits and radios: Nominal, all charging. One internal zipper tag was clipped, and we found it greatly improved zipper operation. I will apply tag removal to remaining suits tomorrow.

Summary of GreenHab operations: Crops watered at 5pm, Timer for supplemental lighting was prepared to operate from 6-7 pm. At time of report preparation, lights had come on.

WATER USE: 3 gallons

Heater: On

Supplemental light: On

Harvest: None

Summary of Science Dome operations: Guillaume has been harvesting Spirulina this afternoon, and working with Roger on molds. This afternoon Roger set the molds in the oven for drying.

Dual split: Heat or AC, Of

Summary of RAM operations: (Tools used, work done) The RAM was used as a large airlock for this morning’s EVA to great success. EVA team donned radios and suits in the downstairs of the Hab, where we completed radio checks. They then carried their helmets to the RAM. We mounted their helmets, and the team was prepared for departure. Team arrived at end of EVA with a few minutes to spare, and after re-pressurization, removed their helmets, and returned to the Hab to stow the gear.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Immediately after entering the airlock, Liz began experiencing severe shoulder pain. The team reacted quickly to release her suit pack straps and support its weight temporarily relieving her immediate issue. Following depressurization, Liz was extricated from her helmet, and encouraged to return to the Hab and treat her muscle pain. When the remainder of the team returned, we found her somewhat recovered, and in much better spirits.

Questions, concerns and requests to Mission Support:

Liz reported that while the adjustment of the lower strap on her suit did provide a noticeable improvement in the muscles of the lower back, clearly additional work will be needed to realize relief for the shoulder area as well for individuals sharing her stature.

Copyright © The Mars Society. All rights reserved. | Main Site